OpenStudy (anonymous):

Jack usually mows his lawn in 6 hrs. Marilyn can mow the same yard in 4 hrs. How much time would it take for them to mow the lawn together?

OpenStudy (anonymous):

No, way less.

OpenStudy (anonymous):

oh wait, its 2.5 hours!

OpenStudy (anonymous):

No. It's 2.4 hours.

OpenStudy (anonymous):

>.< lol really?

OpenStudy (anonymous):

Yes. 1/6 + 1/4 is 1/2.4 So the answer is 2.4 hrs.

OpenStudy (anonymous):

So 1/4+1/6=1/x add 1/4 and 1/6 and do 1/that value to get 2.4

OpenStudy (anonymous):

ah yes, read wrong on my calc

OpenStudy (anonymous):

ok, Thanks everyone

OpenStudy (anonymous):

(12/5) hrs =2.4 hrs